# Data-Driven Parametrization of Molecular Mechanics Force Fields for Expansive Chemical Space Coverage

### v1

@@ 3.3 Equilibrium and Off-Equilibrium Conformations
In addition to predict the torsional PES of molecules, the OpenFFBenchmark45 dataset is used to benchmark the performance of various force fields in terms of predicting unconstrained equilibrium conformations and associated energies.
Figure 5: Histograms of different metrics on OpenFFBenchmark. The accuracy of force field-relaxed geometry relative to QM-relaxed references is quantified with (a) RMSD and (b) Torsion Fingerprint Deviation (TFD) scores. The energetic accuracy is quantified by (c) $\Delta\Delta E$ distributions. All benchmark results for “OPLS4 cst”, and the detailed protocols to calculate the benchmark results are obtained from ref 45.
As shown in Fig. 5 (a) and (b), the distribution of ByteFF-gopt and ByteFF-joint in terms of RMSD and TFD are both concentrated near zero, with higher peak values near zero and narrower distributions than other force fields, demonstrating superior consistencies with QM-optimized molecular conformations. Particularly, both ByteFF-gopt and ByteFF-joint significantly outperform the “OPLS4 cst” model (OPLS4 with refined parameters using the FFBuilder tool for each molecule individually), which is considered the state-of-the-art MM force field in current reports. Additionally, we calculated the relative energy differences ($\Delta\Delta E$) to assess energetic agreement between force field-optimized conformations and the QM-optimized ones (Fig. 5 (c)). Both ByteFF-gopt and ByteFF-joint exhibit sharp peaks around zero in their $\Delta\Delta E$ distribution, significantly outperform all other force fields, indicating that both of them reproduce correctly the QM relative energies between conformers. The accurate predictions of equilibrium conformations ensures the correct conformations being sampled in MD simulations, while the precise prediction of $\Delta\Delta E$ guarantees the Boltzmann distribution in the conformational space is properly reproduced.
@@
In addition to the equilibrium conformations, it is equally important for the MD simulations to accurately predict the off-equilibrium conformations. We used several datasets to comprehensively assess the prediction accuracy of both energy and force predictions from various force fields, quantified by RMSE relative to QM references (Table 2). Among all five subsets, ByteFF-gopt achieved higher accuracy than both GAFF-2.2 and OpenFF-2.0, but less accurate than Espaloma-0.3.0. Importantly, ByteFF-gopt achieved this accuracy with relaxed geometry, energy, and partial hessian, without using any force labels in the training process. With additional off-equilibrium energy and forces training data, we further trained ByteFF-joint to improve the force prediction capabilities of ByteFF-gopt. Remarkably, after fine-tuning with just around 10k data in the off-equilibrium dataset—a significantly smaller dataset than that used by Espaloma-0.3.0—the performance of ByteFF-joint surpassed that of its competitors. Rigorous procedures were followed in the benchmarking of ByteFF models, which ensured no data leakage between the training and testing data.
As a result, ByteFF-joint achieves state-of-the-art performance in almost all subsets, except for the SPICE-Pubchem dataset, where ByteFF-joint shows marginally higher force RMSE compared to Espaloma-0.3.0. As discussed earlier, it is inherently challenging of an MMFF to perfectly replicate the QM PES. There is a trade-off between predicting the PES of relaxed-geometry and off-equilibrium conformations, and we focused more on the relaxed-geometry performances during the training process, since this is more relevant to the sampling of local minima in molecular dynamics simulations and is more stressed in the latest large-scale collaborative industrial benchmark45. Despite this, ByteFF-joint achieves overall state-of-the-art performance in all benchmarks considered in this study.
@@
SPICE-Pubchem Energy 4.8 4.4 2.3 3.4 2.2
Force 14.1 14.0 6.5 9.9 6.9
@@
Force 10.7 12.5 6.0 8.0 5.5
SPICE-Dipeptide Energy 5.2 4.5 3.3 3.8 2.3
Force 12.5 12.4 8.1 9.6 5.5
RNA-Diverse Energy 6.7 5.4 3.9 6.3 3.5
Force 15.8 18.4 4.5 11.1 3.8
RNA-Trinucleotide Energy 6.2 6.0 3.8 6.0 3.4
Force 16.5 18.7 4.4 11.3 3.7
Table 2: Comparison of energy and force RMSE for different dataset and force fields. SPICE49, RNA, peptide.
@@ 3.4 Limitations
Despite the excellent performance of ByteFF compared to other force fields, several limitations still exist. As an MMFF with a minimal analytical functional, ByteFF inherently lacks the capability of modeling more complicated effects in intramolecular PES, such as bond-torsion and angle-torsion couplings. Moreover, while intramolecular parameters are carefully trained against QM references, non-bonded parameters are directly inherited from GAFF to ensure Amber compatibility, leaving room for further refinement. Addressing these limitations presents a path forward for the further improvements of ByteFF.
@@ 4 Conclusion
In this work, we developed ByteFF, an Amber-compatible MMFF with broad chemical space coverage and state-of-the-art accuracy. Through Morgan-based torsional fingerprints and t-SNE analysis, we demonstrated that our dataset covers superior chemical space. Based on this strong data foundation, ByteFF leverages an edge-augmented, symmetry-preserving GNN model to predict MMFF parameters. We also carefully designed our training strategy, enhanced by techniques such as partial Hessian fitting, an iterative optimization-and-training scheme, and effective model ensembling for prediction improvements and uncertainty quantification. Benefitting from the comprehensive datasets and exquisitely designed training strategy, ByteFF outperforms existing MM force fields across multiple benchmarks, delivering accurate predictions of torsional energy profiles, equilibrium conformations, and off-equilibrium energies and forces. We believe ByteFF will significantly benefit biomolecular simulations for various purposes.

### v2

@@ 3.3 Equilibrium and Off-Equilibrium Conformations
In addition to predict the torsional PES of molecules, the OpenFFBenchmark51 dataset is used to benchmark the performance of various force fields in terms of predicting unconstrained equilibrium conformations and associated energies.
Figure 5: Histograms of different metrics on OpenFFBenchmark dataset. The accuracy of force field-relaxed geometry relative to QM-relaxed references is quantified with (a) RMSD and (b) Torsion Fingerprint Deviation (TFD) scores. The energetic accuracy is quantified by (c) $\Delta\Delta E$ distributions. All benchmark results for “OPLS4 cst”, and the detailed protocols to calculate the benchmark results are obtained from ref 51.
As shown in Fig. 5 (a) and (b), the distribution of ByteFF-gopt and ByteFF-joint in terms of RMSD and TFD are both concentrated near zero, with higher peak values near zero and narrower distributions than other force fields, demonstrating superior consistencies with QM-optimized molecular conformations. Particularly, both ByteFF-gopt and ByteFF-joint significantly outperform the “OPLS4 cst” model (OPLS4 with refined parameters using the FFBuilder tool for each molecule individually), which is considered the state-of-the-art MMFF in current reports. Additionally, we calculated the relative energy differences ($\Delta\Delta E$) to assess energetic agreement between force field-optimized conformations and the QM-optimized ones (Fig. 5 (c)). Both ByteFF-gopt and ByteFF-joint exhibit sharp peaks around zero in their $\Delta\Delta E$ distribution, significantly outperform all other force fields, indicating that both of them reproduce correctly the QM relative energies between conformers. The accurate predictions of equilibrium conformations ensures the correct conformations being sampled in MD simulations, while the precise prediction of $\Delta\Delta E$ guarantees the Boltzmann distribution in the conformational space is properly reproduced.
@@
In addition to the equilibrium conformations, it is equally important for the MD simulations to accurately predict the off-equilibrium conformations. We used several datasets to comprehensively assess the prediction accuracy of both energy and force predictions from various force fields, quantified by RMSE relative to QM references (Table 2). Among all five subsets, ByteFF-gopt achieved higher accuracy than both GAFF-2.2 and OpenFF-2.0, but less accurate than Espaloma-0.3.0. Importantly, ByteFF-gopt achieved this accuracy with relaxed geometry, energy, and partial hessian, without using any force labels in the training process. With additional off-equilibrium energy and forces training data, we further trained ByteFF-joint to improve the force prediction capabilities of ByteFF-gopt. Remarkably, after fine-tuning with just around 10k data in the off-equilibrium dataset—a significantly smaller dataset than that used by Espaloma-0.3.0—the performance of ByteFF-joint surpassed that of its competitors.
As a result, ByteFF-joint achieves state-of-the-art performance in almost all subsets, except for the SPICE-Pubchem dataset, where ByteFF-joint shows marginally higher force RMSE compared to Espaloma-0.3.0. As discussed earlier, it is inherently challenging of an MMFF to perfectly replicate the QM PES. There is a trade-off between predicting the PES of relaxed-geometry and off-equilibrium conformations, and we focused more on the relaxed-geometry performances during the training process, since this is more relevant to the sampling of local minima in molecular dynamics simulations and is more stressed in the latest large-scale collaborative industrial benchmark51. Despite this, ByteFF-joint achieves overall state-of-the-art performance in all benchmarks considered in this study.
Table 2: Comparison of energy RMSE (in kcal/mol) and force RMSE (in kcal/mol/Å) for different datasets and force fields.
@@
SPICE-Pubchem Energy 4.8 4.4 2.3 3.4 2.3
Force 14.1 14.0 6.5 9.9 6.7
@@
Force 10.7 12.5 6.0 8.0 5.4
SPICE-Dipeptide Energy 5.2 4.5 3.3 4.0 2.3
Force 12.5 12.4 8.1 9.9 5.3
RNA-Diverse Energy 6.7 5.4 3.9 4.6 3.5
Force 15.8 18.4 4.5 9.7 3.7
RNA-Trinucleotide Energy 6.2 6.0 3.8 4.4 3.4
Force 16.5 18.7 4.4 10.0 3.6
@@ 3.4 Limitations
Despite the excellent performance of ByteFF compared to other force fields, several limitations still exist. As an MMFF with a minimal analytical functional, ByteFF inherently lacks the capability of modeling more complicated effects in intramolecular PES, such as bond-torsion and angle-torsion couplings. Moreover, while intramolecular parameters are carefully trained against QM references, non-bonded parameters are directly inherited from GAFF-2.2 to ensure Amber compatibility, leaving room for further refinement. Addressing these limitations presents a path forward for the further improvements of ByteFF.
@@ 4 Conclusion
In this work, we developed ByteFF, an Amber-compatible MMFF with broad chemical space coverage and state-of-the-art accuracy. Through Morgan-based torsional fingerprints and t-SNE analysis, we demonstrated that our dataset covers superior chemical space. Based on this strong data foundation, ByteFF leverages an edge-augmented, symmetry-preserving GNN model to predict MMFF parameters. We also carefully designed our training strategy, enhanced by techniques such as partial Hessian fitting, an iterative optimization-and-training scheme, and effective model ensembling for prediction improvements and uncertainty quantification. Benefitting from the comprehensive datasets and exquisitely designed training strategy, ByteFF outperforms existing MMFFs across multiple benchmarks, delivering accurate predictions of torsional energy profiles, equilibrium conformations, and off-equilibrium energies and forces. We believe ByteFF will significantly benefit biomolecular simulations for various purposes.